OpenStudy (anonymous):

Solve the equation for Y. 8x-9y=11 A) y=-9/8x - 9/11 B) y=9/8x+11/8 C) y=-8x-11 D) y=8/9x-11/9 plz help will medal!+11 more questions

OpenStudy (anonymous):

@Peach-es1234 wow thx for your help can you help me with 11 more? if you want i can open up a question so that i can give you more medals

OpenStudy (anonymous):

I will try :)

OpenStudy (anonymous):

um i got the answer 2.5 thats not one of the options im confused ;-;

OpenStudy (anonymous):

do you want me to open up a new question so you can get more medals?

OpenStudy (anonymous):

Sure if you want too.

OpenStudy (anonymous):

Simplifying 8x + -9y = 11 Solving 8x + -9y = 11 Move all terms containing Y to the left, all other terms to the right. Add '-8x' to each side of the equation. 8x + -8x + -9y = 11 + -8x Combine like terms: 8x + -8x = 0 0 + -9y = 11 + -8x -9y = 11 + -8x Add '9y' to each side of the equation. -9y + 9y = 11 + -8x + 9y Combine like terms: -9y + 9y = 0 0 = 11 + -8x + 9y Simplifying 0 = 11 + -8x + 9y Combine like terms: 0 + -11 = -11 -11 + 8x + -9y = 11 + -8x + 9y + -11 + 8x + -9y Reorder the terms: -11 + 8x + -9y = 11 + -11 + -8x + 8x + 9y + -9y Combine like terms: 11 + -11 = 0 -11 + 8x + -9y = 0 + -8x + 8x + 9y + -9y -11 + 8x + -9y = -8x + 8x + 9y + -9y Combine like terms: -8x + 8x = 0 -11 + 8x + -9y = 0 + 9y + -9y -11 + 8x + -9y = 9y + -9y Combine like terms: 9y + -9y = 0 -11 + 8x + -9y = 0 I think the answer is A.

OpenStudy (anonymous):

@Peach-es1234 thx let me open up a new question

OpenStudy (anonymous):

Was it correct?

OpenStudy (anonymous):

i dont know i think it was

OpenStudy (anonymous):

Ok, :)